# New Haven, United States

### v1

@@ 1 Introduction
While standard topic models can identify general themes, they often lack specificity and coherence in domain-focused applications. To address this, guided topic modeling approaches [6, 4] allow users to introduce domain-specific topic seeds, improving the model’s alignment with expert knowledge and generating topics that are more coherent and more relevant to specific applications. For instance, models such as SeededLDA [6] and CorEx [4] leverage seed word lists to help align the process of topic discovery, enhancing the interpretability and precision of topics based on the user’s goals. However, even with guided approaches, these models still face limitations in capturing nuanced relationships and require humans in the loop to guide the topic discovery, which is time-consuming and labor-intensive.
@@
To address these challenges, we propose an LLM-assisted Iterative Topic Augmentation framework (LITA), designed to balance LLM utility with cost-effectiveness. Our framework begins with user-provided topic seeds and employs an embedding model to cluster documents, followed by an iterative refinement and augmentation process, making it distinct from conventional methods. Specifically, LITA identifies ambiguous documents based on their proximity to cluster centroids and employ an LLM to assign only these documents to existing or new topics. By reassigning and dynamically augmenting new topics as needed, LITA achieves high-quality, coherent topics while minimizing LLM queries, enabling more scalable and adaptable topic modeling.
@@ 2 Related Works
Topic modeling is a text clustering technique for uncovering thematic structures within a corpus. One notable foundational algorithm, LDA [1], infers document-topic and topic-word distributions, providing a framework for users to explore prevalent topics. Building on this, guided topic modeling allows users to inject domain knowledge into the topic identification process, tailoring results to specific needs and generating more coherent topics. For instance, SeededLDA [6] and CorEx [4] enable users to provide topic seed sets, helping shape the model’s topic structure in line with their desired themes. Afterward, with the development of pre-trained language models such as BERT [12], frameworks like BERTopic [5] have leveraged contextualized embeddings for topic clustering by applying traditional clustering algorithms, such as HDBSCAN or K-means, to these embeddings. Our proposed framework is inspired by BERTopic but introduces a significant advancement: it starts with human-provided topic seeds and then integrates the LLM’s text understanding capability, iteratively refining and augmenting clustering results based on LLM feedback.
@@ 4.2 Results & Analysis
Figure 2: Main Results: we report LITA’s performance across iterations (e.g., from 1 to 5 on CLINC(D) dataset) to show how the performance changes. Although other baselines do not implement an iterative scheme, we report their performances on different numbers of topics ($K$) that match the number of topics generated by LITA in different iterations to ensure a fair comparison.
@@
Fig 2 (a) examines topic coherence (measured by NPMI) across multiple baselines. The findings indicate that LITA consistently demonstrates superior performance in terms of topic coherence. It achieves the highest NPMI values across most cluster configurations. In comparison, PromptTopic, BERTopic and CorEX exhibit similar but slightly lower coherence levels. Traditional methods such as LDA and SeededLDA perform substantially worse, whereas SeededLDA produces negative NPMI scores, highlighting its inability to generate semantically coherent topics. In addition to coherence, Fig 2 (b) shows that LITA demonstrates competitive performance in topic diversity. To maintain high diversity while preserving coherence is a critical attribute for practical applications. While SeededLDA records a maximum diversity score in the 20 Newsgroups, its negative coherence metrics indicate a tradeoff that undermines semantic quality. The comparison results for NMI and cluster accuracy are presented in Fig.2 (c) and (d), respectively. The results reveal an upward trend in NMI and accuracy for LITA during iterations (as the number of clusters increases), reflecting improved alignment with the ground truth as granularity becomes finer. Although PromptTopic achieves results comparable to LITA in NMI and accuracy, LITA’s enhanced efficiency offers a clear advantage, which will be explored in detail in subsequent sections.
@@
Previous prompt-based methods necessitate a greater number of LLM requests due to their workflow. For instance, PromptTopic processes each document individually to assign topics, resulting in 7,532 and 4,500 API calls for the 20 Newsgroups and CLINC(D) datasets, respectively. Furthermore, PromptTopic requires additional API access during the topic collapse phase. In contrast, LITA strategically limits LLM invocations to only the evaluation of ambiguous instances. As illustrated in Figure 3, the number of ambiguous instances decreases substantially across iterations for both datasets, reaching topic convergence within just 5–7 iterations. As a result, LITA requires just 1,325 and 487 API calls on the same datasets. Compared to PromptTopic, LITA reduces over 80% LLM requests on both 20 Newsgroups and CLINC(D) datasets. Considering to LITA’s competitive performances reported in Fig. 2, LITA identifies topics of better quality while achieving computational efficiency, making LITA as cost-effective and scalable solution for real-world applications.
@@ 5 Conclusion
This paper presents LITA, an LLM-assisted iterative topic augmentation framework that balances cost-effectiveness and high-quality topic modeling. By leveraging LLMs to refine ambiguous documents and expand topics dynamically, LITA outperforms baseline methods across key metrics, including topic coherence, diversity, NMI, and clustering accuracy. Its efficient integration of LLM feedback without over-relying on costly API queries distinguishes it from other approaches. While LITA shows strong empirical results, limitations such as dependency on user-provided seeds and dataset-specific performance variations remain. Future work could focus on expanding to larger datasets. In conclusion, LITA provides a scalable, efficient solution for advancing topic modeling and demonstrates the potential of LLMs in this domain.

### v2

@@ 1 Introduction
While standard topic models can identify general themes, they often lack specificity and coherence in domain-focused applications. To address this, guided topic modeling approaches [7, 5, 11, 3] allow users to introduce domain-specific topic seeds, improving the model’s alignment with expert knowledge and generating topics that are more coherent and more relevant to specific applications. For instance, models such as SeededLDA [7], CorEx [5], and GTM [3] leverage seed word lists to help align the process of topic discovery, enhancing the interpretability and precision of topics based on the user’s goals. However, even with guided approaches, these models still face limitations in capturing nuanced relationships and require humans in the loop to guide the topic discovery, which is time-consuming and labor-intensive.
@@
To address these challenges, we propose an LLM-assisted Iterative Topic Augmentation framework (LITA) [^1] Our codes are available at https://github.com/Text-Analytics-and-Retrieval/LITA-LLM-Iterative-Topic-Augmentation., designed to balance LLM utility with cost-effectiveness. Our framework begins with user-provided topic seeds and employs an embedding model to cluster documents, followed by an iterative refinement and augmentation process, making it distinct from conventional methods. Specifically, LITA identifies ambiguous documents based on their proximity to cluster centroids and employ an LLM to assign only these documents to existing or new topics. By reassigning and dynamically augmenting new topics as needed, LITA achieves high-quality, coherent topics while minimizing LLM queries, enabling more scalable and adaptable topic modeling.
@@ 2 Related Works
Topic modeling is a text clustering technique for uncovering thematic structures within a corpus. One notable foundational algorithm, LDA [1], infers document-topic and topic-word distributions, providing a framework for users to explore prevalent topics. Building on this, guided topic modeling allows users to inject domain knowledge into the topic identification process, tailoring results to specific needs and generating more coherent topics. SeededLDA [7] is the pioneer that leverages seed word lists to help align the process of topic discovery. CorEx [5] takes user-provided guidance as anchor words in the embedding space and learns maximally informative topics. CatE [11] simultaneously learns the embedding for user-provided category names, documents, and words, resulting in a category-aligned topic embedding model. GTM [3] modifies the LDA generative process to incorporate user-provided seed words, incrementally improving the seed words and identifying new topics. These models may face limitations in capturing nuanced relationships because they still learn topic structures from word co-occurrences without using the contextualized embedding captured by advanced language models. They also require humans in the loop to guide the topic discovery, which is time-consuming and labor-intensive. With the development of pre-trained language models such as BERT [14], frameworks like BERTopic [6] have leveraged contextualized embeddings for topic clustering by applying traditional clustering algorithms, such as HDBSCAN or K-means, to these embeddings. Our proposed framework is inspired by BERTopic but incorporates the benefit from guided topic modeling approaches: it starts with human-provided topic seeds and then integrates the LLM’s text understanding capability, iteratively refining and augmenting clustering results based on LLM feedback.
@@ 4.2 Results & Analysis
Table 1: Main Results: we report LITA’s performance across iterations (i.e., 1, 3, 5, and 7 on 20Newsgroups and from 1 to 5 on CLINC(D) dataset). Although other baselines do not implement an iterative scheme, we report their performances on different numbers of topics ($K$) that match the number of topics generated by LITA in different iterations to ensure a fair comparison.
20Newsgroups CLINC(D)
Metric Method 1(5) 3(19) 5(25) 7(28) 1(3) 2(8) 3(11) 4(13) 5(14)
NPMI LITA 0.121 0.277 0.279 0.304 0.132 0.264 0.280 0.312 0.313
PromptTopic 0.126 0.218 0.233 0.254 0.135 0.279 0.289 0.292 0.301
BERTopic 0.120 0.226 0.244 0.259 0.126 0.224 0.249 0.268 0.277
corEx 0.111 0.218 0.249 0.250 0.139 0.243 0.272 0.276 0.282
SeededLDA -0.129 -0.122 -0.112 -0.109 -0.131 -0.157 -0.152 -0.152 -0.149
LDA 0.112 0.198 0.217 0.231 0.117 0.224 0.226 0.240 0.243
Topic Diversity LITA 0.481 0.548 0.542 0.559 0.701 0.665 0.678 0.676 0.689
PromptTopic 0.521 0.527 0.540 0.536 0.672 0.675 0.661 0.649 0.655
BERTopic 0.249 0.309 0.328 0.329 0.526 0.498 0.514 0.528 0.532
corEx 0.431 0.425 0.433 0.445 0.581 0.552 0.549 0.542 0.536
SeededLDA 0.840 0.622 0.633 0.661 0.643 0.561 0.583 0.579 0.564
LDA 0.239 0.246 0.251 0.252 0.344 0.336 0.293 0.299 0.298
NMI LITA 0.457 0.635 0.635 0.637 0.393 0.471 0.531 0.594 0.616
PromptTopic 0.459 0.630 0.632 0.633 0.388 0.467 0.521 0.581 0.612
BERTopic 0.248 0.444 0.460 0.482 0.063 0.455 0.486 0.477 0.496
corEx 0.055 0.184 0.214 0.222 0.105 0.257 0.262 0.288 0.258
SeededLDA 0.371 0.479 0.434 0.454 0.056 0.132 0.174 0.206 0.164
LDA 0.177 0.218 0.225 0.245 0.050 0.096 0.101 0.116 0.102
Accuracy LITA 0.241 0.560 0.549 0.547 0.288 0.516 0.550 0.616 0.618
PromptTopic 0.239 0.555 0.542 0.542 0.287 0.514 0.533 0.601 0.609
BERTopic 0.145 0.320 0.367 0.368 0.140 0.382 0.389 0.396 0.392
corEx 0.099 0.177 0.211 0.188 0.192 0.352 0.345 0.317 0.316
SeededLDA 0.224 0.427 0.398 0.382 0.201 0.268 0.250 0.245 0.219
LDA 0.135 0.228 0.229 0.241 0.168 0.230 0.206 0.238 0.186
@@
Table 1 presents the evaluation results in four metrics. In terms of topic coherence (measured by NPMI), LITA consistently demonstrates superior performance. LITA achieves the highest or second highest NPMI values in most cluster configurations. In comparison, PromptTopic, BERTopic, and CorEX exhibit similar but slightly lower coherence levels. Traditional methods such as LDA and SeededLDA perform substantially worse, whereas SeededLDA produces negative NPMI scores, highlighting its inability to generate semantically coherent topics. In addition to coherence, LITA demonstrates competitive performance in topic diversity. Maintaining high diversity while preserving coherence is a critical attribute for practical applications. Although SeededLDA records a maximum diversity score in the 20Newsgroups, its negative coherence metrics indicate a trade-off that undermines semantic quality. For clustering performance, the results reveal an upward trend in NMI and accuracy for LITA during iterations (as the number of clusters increases), reflecting improved alignment with the ground truth as the granularity becomes finer. Although PromptTopic achieves results comparable to LITA in NMI and accuracy, LITA’s enhanced efficiency offers a clear advantage, which will be explored in detail in the subsequent section.
@@
Previous prompt-based methods necessitate a greater number of LLM requests due to their workflow. For instance, PromptTopic processes each document individually to assign topics, resulting in 7,532 and 4,500 API calls for the 20 Newsgroups and CLINC(D) datasets, respectively. Furthermore, PromptTopic requires additional API access during the topic collapse phase. In contrast, LITA strategically limits LLM invocations to only the evaluation of ambiguous instances. As illustrated in Figure 2, the number of ambiguous instances decreases substantially across iterations for both datasets, reaching topic convergence within just 5–7 iterations. As a result, LITA requires just 1,325 and 487 API calls on the same datasets. Compared to PromptTopic, LITA reduces over 80% LLM requests on both 20 Newsgroups and CLINC(D) datasets. Considering to LITA’s competitive performances reported in Table 1, LITA identifies topics of better quality while achieving computational efficiency, making LITA as cost-effective and scalable solution for real-world applications.
@@ 5 Conclusion
This paper presents LITA, an LLM-assisted iterative topic augmentation framework that balances cost-effectiveness and high-quality topic modeling. By leveraging LLMs to refine ambiguous documents and expand topics dynamically, LITA outperforms baseline methods across key metrics, including topic coherence, diversity, NMI, and clustering accuracy. Its efficient integration of LLM feedback without over-relying on costly API queries distinguishes it from other approaches. While LITA shows strong empirical results, limitations such as dependency on user-provided seeds and dataset-specific performance variations remain. Future work could focus on evaluating LITA on larger-scale and more domain-specific datasets to assess its generalizability and robustness. In conclusion, LITA provides a scalable and efficient solution for advancing topic modeling and demonstrates the potential of LLMs in this domain.